Three identical point masses are placed at the corners of an equilateral triangle of sides"R". If the mass of each particle is "M" then the magnitude of the netforce on each particle is??

1
Expert's answer
2022-03-27T18:20:33-0400

"F=\frac{Gm^2}{(\frac 23\cdot\frac{r\sqrt 3}2)^2}=\frac{3Gm^2}{r^2}."